# Diederich–Fornæss index and global regularity of the complex Green operator: domains with comparable Levi eigenvalues

### v1

@@ 4. The main estimate
(4.30) $\sideset{}{{}^{\prime}}{\sum}_{J}|\alpha_{\rho_{\eta}}(L^{J}_{\mathcal{P}^{-,\eta}u})|^{2}\\ \lesssim(1-\eta)\left(\|\overline{\partial}_{b}(e^{-h_{\eta}/2}\mathcal{P}^{-,\eta}u)\|_{-h_{\eta}}^{2}+\|\overline{\partial}_{b,-h_{\eta}}^{*}(e^{-h_{\eta}/2}\mathcal{P}^{-,\eta}u)\|_{-h_{\eta}}^{2}+\|\overline{\partial}_{b}u\|^{2}+\|\overline{\partial}_{b}^{*}u\|^{2}\right)\\ +C_{\eta}\|u\|_{-1}^{2}\;.\;\;\;\;\;\;\;\;\;$
@@
(4.34) $\sideset{}{{}^{\prime}}{\sum}_{J}|\alpha_{\rho_{\eta}}(L^{J}_{\mathcal{P}^{0,\eta}u})|^{2}\\ \lesssim(1-\eta)\left(\|\overline{\partial}_{b}(e^{-h_{\eta}/2}\mathcal{P}^{0,\eta}u)\|_{-h_{\eta}}^{2}+\|\overline{\partial}_{b,-h_{\eta}}^{*}(e^{-h_{\eta}/2}\mathcal{P}^{0,\eta}u)\|_{-h_{\eta}}^{2}+\|\overline{\partial}_{b}u\|^{2}+\|\overline{\partial}_{b}^{*}u\|^{2}\right)\\ +C_{\eta}\|u\|_{-1}^{2}\;.\;\;\;\;\;\;\;\;\;$
@@
(4.35) $\bar{\partial}^{\star}_{b,-h_{\eta}}(e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u)\\ =e^{-h_{\eta}}\overline{\partial}_{b}^{*}(e^{h_{\eta}}e^{-h_{\eta}/2}\mathcal{P}^{+,\eta})=e^{-h_{\eta}}\overline{\partial}_{b}^{*}(e^{h_{\eta}/2}\mathcal{P}^{+,\eta}u)\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\\ \;\;=e^{-h_{\eta}/2}\bar{\partial}_{b}^{\star}\mathcal{P}^{+,\eta}u+\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{K}e^{-h_{\eta}/2}(L_{j}h_{\eta})(\mathcal{P}^{+,\eta}u_{jK})\overline{\omega}_{K}+\mathcal{O}(|e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u|)\;,$
@@
(4.37) $\Big\|{\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{K}(L_{j}h_{\eta})\mathcal{P}^{+,\eta}u_{jK}}\Big\|^{2}\\ =\Big\|{\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{K}dh(L_{j})\mathcal{P}^{+,\eta}u_{jK}}\Big\|^{2}=\Big\|\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{K}\left(\alpha_{\eta}-\alpha^{\rho}\right)(L_{j})\mathcal{P}^{+,\eta}u_{jK}\Big\|^{2}\\ \lesssim\sideset{}{{}^{\prime}}{\sum}_{K}\|\alpha_{\eta}\big(\sum_{j}(\mathcal{P}^{+,\eta}u)_{jK}L_{j}\big)\|^{2}+\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{K}\|\alpha^{\rho}(L_{j})\mathcal{P}^{+,\eta}u_{jK}\|^{2}\\ \lesssim\sideset{}{{}^{\prime}}{\sum}_{K}\|\alpha_{\eta}(L_{\mathcal{P}^{+,\eta}u}^{K})\|^{2}+\left\|u\right\|^{2}.$
We have used that $\|\mathcal{P}^{+,\eta}u\|^{2}\leq\|u\|^{2}$. For the last inequality, see the discussion after the definition of $L_{u}^{J}$ immediately before (3.8): $\sum_{j}(\mathcal{P}^{+,\eta}u)_{jK}L_{j}$ is the $L^{K}$–term’ for $\mathcal{P}^{+,\eta}$ in the boundary chart, and we may replace the sum over $K$ by the corresponding sum for the ‘actual’ $L_{\mathcal{P}^{+,\eta}u}^{K}$–terms (i.e. the ones formed in the Euclidean chart). (4.36) and (4.37) combine to give
(4.38) $\|\bar{\partial}_{b,-h_{\eta}}^{\star}(e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u)\|_{-h_{\eta}}^{2}\lesssim\left\|\bar{\partial}_{b}^{\star}\mathcal{P}^{+,\eta}u\right\|^{2}+\sideset{}{{}^{\prime}}{\sum}_{K}\left\|\alpha_{\eta}(L_{\mathcal{P}^{+,\eta}u}^{K})\right\|^{2}+\left\|u\right\|^{2}.$
@@
(4.39) $\begin{split}\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}\left|\alpha_{\eta}(L_{\mathcal{P}^{+,\eta}u}^{K})\right|^{2}&\lesssim(1-\eta)\Big(\|\bar{\partial}_{b}(e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u)\|_{-h_{\eta}}^{2}+\left\|\bar{\partial}_{b}^{\star}\mathcal{P}^{+,\eta}u\right\|^{2}+\left\|u\right\|^{2}\\ &\qquad+\sideset{}{{}^{\prime}}{\sum}_{K}\left\|\alpha_{\eta}(L_{\mathcal{P}^{+,\eta}u}^{K})\right\|^{2}+\left\|\bar{\partial}_{b}u\right\|^{2}+\|\overline{\partial}_{b}^{*}u\|^{2}\Big)+C_{\eta}\left\|u\right\|_{-1}^{2}.\end{split}$
For $\eta$ sufficiently close to $1$, we absorb $(1-\eta)\sum^{{}^{\prime}}_{K}\left\|\alpha_{\eta}(L_{\mathcal{P}^{+,\eta}u}^{K})\right\|^{2}=(1-\eta)\sum_{K}^{\prime}\int_{b\Omega}\left|\alpha_{\eta}(L_{\mathcal{P}^{+,\eta}u}^{K})\right|^{2}$ into the left hand side of (4.39) and use $\left\|u\right\|^{2}\lesssim\left\|\bar{\partial}_{b}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}u\right\|^{2}$ to obtain
(4.40) $\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}\left|\alpha_{\eta}(L_{\mathcal{P}^{+,\eta}u}^{K})\right|^{2}\lesssim(1-\eta)\Big(\|\bar{\partial}_{b}(e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u)\|_{-h_{\eta}}^{2}+\|\overline{\partial}_{b}^{*}\mathcal{P}^{+,\eta}u\|^{2}\\ +\left\|\bar{\partial}_{b}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}u\right\|^{2}\Big)+C_{\eta}\left\|u\right\|_{-1}^{2}.$
@@
The last inequality follows as in (4.37). For $J$ fixed, let $J^{\prime}\subset J$ be a (q1)(q-1)-tuple that does not contain $j$. Then $\mathcal{P}^{+,\eta}(u_{J}\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}})=(\mathcal{P}^{+,\eta}u_{J})\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}}$, and $\sum_{s}(\mathcal{P}^{+,\eta}u_{J}\overline{w_{j}}\wedge\overline{\omega_{J^{\prime}}})_{sJ^{\prime}}L_{s}=\mathcal{P}^{+,\eta}u_{J}L_{j}$ (the sum is only over $s\notin J^{\prime}$, which forces $s=j$ for a nonvanishing term). That is, $\mathcal{P}^{+,\eta}u_{J}L_{j}=L^{J^{\prime}}_{\mathcal{P}^{+,\eta}(u_{J}\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}})}$. We now apply (4.40) with $u$ replaced by $u_{J}\overline{\omega_{j}}\wedge\overline{\omega_{J}}$. Then
@@
(4.43) $\int_{b\Omega}\left|\alpha_{\eta}(\mathcal{P}^{+,\eta}u_{J}L_{j})\right|^{2}=\int_{b\Omega}\left|\alpha_{\eta}\big(L^{J^{\prime}}_{\mathcal{P}^{+,\eta}(u_{J}\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}})}\big)\right|^{2}\leq\sum_{K}\int_{b\Omega}\left|\alpha_{\eta}\big(L^{K}_{\mathcal{P}^{+,\eta}(u_{J}\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}})}\big)\right|^{2}\\ \lesssim(1-\eta)\Big(\|\bar{\partial}_{b}(e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\|_{-h_{\eta}}^{2}+\left\|\bar{\partial}_{b}^{\star}(\mathcal{P}^{+,\eta}u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\\ +\left\|\bar{\partial}_{b}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\Big)+C_{\eta}\left\|u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}}\right\|_{-1}^{2}.$
The last inequality follows from (4.40) (and again the discussion after the definition of $L_{u}^{J}$).
@@
Next, we estimate the first term on the right hand side of (4.43).
@@
(4.44) $\|\bar{\partial}_{b}(e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\|_{-h_{\eta}}^{2}\lesssim\sum_{s}\|\overline{L}_{s}(e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u_{J})\|_{-h_{\eta}}^{2}+\|e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u_{J}\|_{-h_{\eta}}^{2}\\ \lesssim\sum_{s}\left(\left\|(\overline{L}_{s}h_{\eta})\mathcal{P}^{+,\eta}u_{J}\right\|^{2}+\left\|\overline{L}_{s}\mathcal{P}^{+,\eta}u_{J}\right\|^{2}\right)+\left\|\mathcal{P}^{+,\eta}u_{J}\right\|^{2}\\ \lesssim\sum_{s}\left\|(\overline{L}_{s}h_{\eta})\mathcal{P}^{+,\eta}u_{J}\right\|^{2}+\left\|\bar{\partial}_{b}\mathcal{P}^{+,\eta}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}\mathcal{P}^{+,\eta}u\right\|^{2}+\left\|\mathcal{P}^{+,\eta}u\right\|^{2}\\ \lesssim\sum_{s}\left\|(\overline{L}_{s}h_{\eta})\mathcal{P}^{+,\eta}u_{J}\right\|^{2}+\left\|\bar{\partial}_{b}\mathcal{P}^{+,\eta}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}\mathcal{P}^{+,\eta}u\right\|^{2}\;.$
@@
For the third inequality, we have used maximal estimates. Starting with the second term on the first line of (4.42) and substituting (4.43) into its right hand side, we get
@@
(4.45) $\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{J}\int_{b\Omega}\left|(\overline{L}_{j}h_{\eta})\mathcal{P}^{+,\eta}u_{J}\right|^{2}\\ \lesssim(1-\eta)\Big(\|\bar{\partial}_{b}(e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\|_{-h_{\eta}}^{2}+\left\|\bar{\partial}_{b}^{\star}(\mathcal{P}^{+,\eta}u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\\ +\left\|\bar{\partial}_{b}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\Big)+C_{\eta}\left\|u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}}\right\|_{-1}^{2}+\|\mathcal{P}^{+,\eta}u\|^{2}\;.$
@@
(4.46) $\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{J}\int_{b\Omega}\left|(\overline{L}_{j}h_{\eta})\mathcal{P}^{+,\eta}u_{J}\right|^{2}\\ \lesssim(1-\eta)\Big(\sum_{s}\left\|(\overline{L}_{s}h_{\eta})\mathcal{P}^{+,\eta}u_{J}\right\|^{2}+\left\|\bar{\partial}_{b}\mathcal{P}^{+,\eta}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}\mathcal{P}^{+,\eta}u\right\|^{2}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\left\|\bar{\partial}_{b}^{\star}(\mathcal{P}^{+,\eta}u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}+\left\|\bar{\partial}_{b}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\Big)\\ +C_{\eta}\left\|u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}}\right\|_{-1}^{2}+\|\mathcal{P}^{+,\eta}u\|^{2}\;.$
@@
Choosing $(1-\eta)$ small enough and absorbing $(1-\eta)(\sum_{s}\left\|(\overline{L}_{s}h_{\eta})u_{J}\right\|^{2})$ into the left hand side gives
(4.47) $\sum_{j}\left\|(\overline{L}_{j}h_{\eta})\mathcal{P}^{+,\eta}u_{J}\right\|^{2}\\ \lesssim(1-\eta)\Big(\left\|\bar{\partial}_{b}\mathcal{P}^{+,\eta}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}\mathcal{P}^{+,\eta}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}(\mathcal{P}^{+,\eta}u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\\ \;\;\;\;\;\;\;\;\;+\left\|\bar{\partial}_{b}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\Big)+\|u\|^{2}+C_{\eta}\|u\|_{-1}^{2}\;.$
We have used that $\|\mathcal{P}^{+,\eta}u\|^{2}\lesssim\|u\|^{2}$ and $\left\|u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}}\right\|_{-1}^{2}\lesssim\|u\|_{-1}^{2}$.
Starting with the first inequality in (4.42) and using (4.47) gives us
(4.48) $\left\|\bar{\partial}_{b}h_{\eta}\wedge\mathcal{P}^{+,\eta}u\right\|^{2}\\ \lesssim(1-\eta)\Big(\left\|\bar{\partial}_{b}\mathcal{P}^{+,\eta}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}\mathcal{P}^{+,\eta}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}(\mathcal{P}^{+,\eta}u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\\ \;\;\;\;\;\;\;\;\;+\left\|\bar{\partial}_{b}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\Big)+\|u\|^{2}+C_{\eta}\|u\|_{-1}^{2}\;.$
Combining (4.40), (4.41), and (4.48), and using $(1-\eta)^{2}\leq(1-\eta)$ and $\left\|u\right\|^{2}\lesssim\left\|\bar{\partial}_{b}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}u\right\|^{2}$, we get

### v2

@@ 4. The main estimate
(4.29) $\sideset{}{{}^{\prime}}{\sum}_{J}|\alpha_{\eta}(\tilde{L}^{J}_{\mathcal{P}^{-,\eta}u})|^{2}\\ \lesssim(1-\eta)\left(\|\overline{\partial}_{b}(e^{-h_{\eta}/2}\mathcal{P}^{-,\eta}u)\|_{-h_{\eta}}^{2}+\|\overline{\partial}_{b,-h_{\eta}}^{*}(e^{-h_{\eta}/2}\mathcal{P}^{-,\eta}u)\|_{-h_{\eta}}^{2}+\|\overline{\partial}_{b}u\|^{2}+\|\overline{\partial}_{b}^{*}u\|^{2}\right)\\ +C_{\eta}\|u\|_{-1}^{2}\;.\;\;\;\;\;\;\;\;\;$
@@
(4.33) $\sideset{}{{}^{\prime}}{\sum}_{J}|\alpha_{\eta}(\tilde{L}^{J}_{\mathcal{P}^{0,\eta}u})|^{2}\\ \lesssim(1-\eta)\left(\|\overline{\partial}_{b}(e^{-h_{\eta}/2}\mathcal{P}^{0,\eta}u)\|_{-h_{\eta}}^{2}+\|\overline{\partial}_{b,-h_{\eta}}^{*}(e^{-h_{\eta}/2}\mathcal{P}^{0,\eta}u)\|_{-h_{\eta}}^{2}+\|\overline{\partial}_{b}u\|^{2}+\|\overline{\partial}_{b}^{*}u\|^{2}\right)\\ +C_{\eta}\|u\|_{-1}^{2}\;.\;\;\;\;\;\;\;\;\;$
@@
(4.34) $\bar{\partial}^{\star}_{b,-h_{\eta}}(e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u)\\ =e^{-h_{\eta}}\overline{\partial}_{b}^{*}(e^{h_{\eta}}e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u)=e^{-h_{\eta}}\overline{\partial}_{b}^{*}(e^{h_{\eta}/2}\mathcal{P}^{+,\eta}u)\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\\ \;\;=e^{-h_{\eta}/2}\bar{\partial}_{b}^{\star}\mathcal{P}^{+,\eta}u+\frac{1}{2}\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{K}e^{-h_{\eta}/2}(L_{j}h_{\eta})(\mathcal{P}^{+,\eta}u_{jK})\overline{\omega}_{K}+\mathcal{O}(|e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u|)\;,$
@@
(4.36) $\Big\|{\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{K}(L_{j}h_{\eta})\mathcal{P}^{+,\eta}u_{jK}}\Big\|^{2}\\ =\Big\|{\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{K}dh(L_{j})\mathcal{P}^{+,\eta}u_{jK}}\Big\|^{2}=\Big\|\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{K}\left(\alpha_{\eta}-\alpha^{\rho}\right)(L_{j})\mathcal{P}^{+,\eta}u_{jK}\Big\|^{2}\\ \lesssim\sideset{}{{}^{\prime}}{\sum}_{K}\|\alpha_{\eta}\big(\sum_{j}(\mathcal{P}^{+,\eta}u)_{jK}L_{j}\big)\|^{2}+\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{K}\|\alpha^{\rho}(L_{j})\mathcal{P}^{+,\eta}u_{jK}\|^{2}\\ \lesssim\sideset{}{{}^{\prime}}{\sum}_{K}\|\alpha_{\eta}(\tilde{L}_{\mathcal{P}^{+,\eta}u}^{K})\|^{2}+\left\|u\right\|^{2}.$
We have used that $\|\mathcal{P}^{+,\eta}u\|^{2}\leq\|u\|^{2}$. Estimates (4.35) and (4.36) combine to give
(4.37) $\|\bar{\partial}_{b,-h_{\eta}}^{\star}(e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u)\|_{-h_{\eta}}^{2}\lesssim\left\|\bar{\partial}_{b}^{\star}\mathcal{P}^{+,\eta}u\right\|^{2}+\sideset{}{{}^{\prime}}{\sum}_{K}\left\|\alpha_{\eta}(\tilde{L}_{\mathcal{P}^{+,\eta}u}^{K})\right\|^{2}+\left\|u\right\|^{2}.$
@@
(4.38) $\begin{split}\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}\left|\alpha_{\eta}(\tilde{L}_{\mathcal{P}^{+,\eta}u}^{K})\right|^{2}&\lesssim(1-\eta)\Big(\|\bar{\partial}_{b}(e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u)\|_{-h_{\eta}}^{2}+\left\|\bar{\partial}_{b}^{\star}\mathcal{P}^{+,\eta}u\right\|^{2}+\left\|u\right\|^{2}\\ &\qquad+\sideset{}{{}^{\prime}}{\sum}_{K}\left\|\alpha_{\eta}(\tilde{L}_{\mathcal{P}^{+,\eta}u}^{K})\right\|^{2}+\left\|\bar{\partial}_{b}u\right\|^{2}+\|\overline{\partial}_{b}^{*}u\|^{2}\Big)+C_{\eta}\left\|u\right\|_{-1}^{2}.\end{split}$
For $\eta$ sufficiently close to $1$, we absorb $(1-\eta)\sum^{{}^{\prime}}_{K}\left\|\alpha_{\eta}(\tilde{L}_{\mathcal{P}^{+,\eta}u}^{K})\right\|^{2}=(1-\eta)\sum_{K}^{\prime}\int_{b\Omega}\left|\alpha_{\eta}(\tilde{L}_{\mathcal{P}^{+,\eta}u}^{K})\right|^{2}$ into the left hand side of (4.38) and use $\left\|u\right\|^{2}\lesssim\left\|\bar{\partial}_{b}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}u\right\|^{2}$ to obtain
(4.39) $\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}\left|\alpha_{\eta}(\tilde{L}_{\mathcal{P}^{+,\eta}u}^{K})\right|^{2}\lesssim(1-\eta)\Big(\|\bar{\partial}_{b}(e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u)\|_{-h_{\eta}}^{2}+\|\overline{\partial}_{b}^{*}\mathcal{P}^{+,\eta}u\|^{2}\\ +\left\|\bar{\partial}_{b}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}u\right\|^{2}\Big)+C_{\eta}\left\|u\right\|_{-1}^{2}.$
@@
The last inequality follows as in (4.36). For each pair $j,J$, let $J^{\prime}\subset J$ be a (q1)(q-1)-tuple that does not contain $j$. Then $\mathcal{P}^{+,\eta}(u_{J}\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}})=(\mathcal{P}^{+,\eta}u_{J})\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}}$, and $\sum_{s}(\mathcal{P}^{+,\eta}u_{J}\overline{w_{j}}\wedge\overline{\omega_{J^{\prime}}})_{sJ^{\prime}}L_{s}=\mathcal{P}^{+,\eta}u_{J}L_{j}$ (the sum is only over $s\notin J^{\prime}$, which forces $s=j$ for a nonvanishing term). That is, $\mathcal{P}^{+,\eta}u_{J}L_{j}=\tilde{L}^{J^{\prime}}_{\mathcal{P}^{+,\eta}(u_{J}\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}})}$. (When $q=1$, $J^{\prime}$ is empty, and $(\mathcal{P}^{+,\eta}u_{J})L_{j}=\tilde{L}_{\mathcal{P}^{+,\eta}}u_{J}\overline{\omega_{j}}$.) We now apply (4.39) with $u$ replaced by $u_{J}\overline{\omega_{j}}\wedge\overline{\omega_{J}}$. Then
@@
(4.42) $\int_{b\Omega}\left|\alpha_{\eta}(\mathcal{P}^{+,\eta}u_{J}L_{j})\right|^{2}=\int_{b\Omega}\left|\alpha_{\eta}\big(\tilde{L}^{J^{\prime}}_{\mathcal{P}^{+,\eta}(u_{J}\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}})}\big)\right|^{2}\leq\sum_{K}\int_{b\Omega}\left|\alpha_{\eta}\big(\tilde{L}^{K}_{\mathcal{P}^{+,\eta}(u_{J}\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}})}\big)\right|^{2}\\ \lesssim(1-\eta)\Big(\|\bar{\partial}_{b}(e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\|_{-h_{\eta}}^{2}+\left\|\bar{\partial}_{b}^{\star}(\mathcal{P}^{+,\eta}u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\\ +\left\|\bar{\partial}_{b}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\Big)+C_{\eta}\left\|u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}}\right\|_{-1}^{2}.$
@@
The last inequality follows from (4.39). Next, we estimate the first term on the right hand side of (4.42).
@@
(4.43) $\|\bar{\partial}_{b}(e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\|_{-h_{\eta}}^{2}\lesssim\sum_{s}\|\overline{L}_{s}(e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u_{J})\|_{-h_{\eta}}^{2}+\|e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u_{J}\|_{-h_{\eta}}^{2}\\ \lesssim\sum_{s}\left(\left\|(\overline{L}_{s}h_{\eta})\mathcal{P}^{+,\eta}u_{J}\right\|^{2}+\left\|\overline{L}_{s}\mathcal{P}^{+,\eta}u_{J}\right\|^{2}\right)+\left\|\mathcal{P}^{+,\eta}u_{J}\right\|^{2}\\ \lesssim\sum_{s}\left\|(\overline{L}_{s}h_{\eta})\mathcal{P}^{+,\eta}u_{J}\right\|^{2}+\left\|\bar{\partial}_{b}\mathcal{P}^{+,\eta}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}\mathcal{P}^{+,\eta}u\right\|^{2}\;.$
@@
For the third inequality, we have used maximal estimates and (2.2). Starting with the second term on the first line of (4.41) and substituting (4.42) into its right hand side, we get
@@
(4.44) $\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{J}\int_{b\Omega}\left|(\overline{L}_{j}h_{\eta})\mathcal{P}^{+,\eta}u_{J}\right|^{2}\\ \lesssim(1-\eta)\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{J}\Big(\|\bar{\partial}_{b}(e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\|_{-h_{\eta}}^{2}+\left\|\bar{\partial}_{b}^{\star}(\mathcal{P}^{+,\eta}u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\\ +\left\|\bar{\partial}_{b}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\Big)\\ +C_{\eta}\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{J}\left\|u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}}\right\|_{-1}^{2}+\|\mathcal{P}^{+,\eta}u\|^{2}\;.$
@@
(4.45) $\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{J}\int_{b\Omega}\left|(\overline{L}_{j}h_{\eta})\mathcal{P}^{+,\eta}u_{J}\right|^{2}\\ \lesssim(1-\eta)\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{J}\Big(\sum_{s}\left\|(\overline{L}_{s}h_{\eta})\mathcal{P}^{+,\eta}u_{J}\right\|^{2}+\left\|\bar{\partial}_{b}\mathcal{P}^{+,\eta}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}\mathcal{P}^{+,\eta}u\right\|^{2}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\left\|\bar{\partial}_{b}^{\star}(\mathcal{P}^{+,\eta}u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}+\left\|\bar{\partial}_{b}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\Big)\\ +C_{\eta}\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{J}\left\|u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}}\right\|_{-1}^{2}+\|\mathcal{P}^{+,\eta}u\|^{2}\;.$
@@
Choosing $(1-\eta)$ small enough and absorbing $(1-\eta)(n-1)\sum_{s}\sum^{{}^{\prime}}_{J}\left\|(\overline{L}_{s}h_{\eta})u_{J}\right\|^{2}$ into the left hand side gives